OpenStudy (canada907cat):

Use the definition of the derivative f'(c) to find f'(c) where f(x)=sqrtx +2x and c=1 I just need to see how the function is set up and I can figure it out from there.

zepdrix (zepdrix):

\[\large\rm f'(c)=\lim_{h\to0}\frac{f(c+h)-f(c)}{h}\]They want us to use c=1,\[\large\rm f'(1)=\lim_{h\to0}\frac{f(1+h)-f(1)}{h}\] \[\large\rm f(\color{orangered}{x})=\sqrt{\color{orangered}{x}}+2\color{orangered}{x}\]At x=1,\[\large\rm f(\color{orangered}{1})=\sqrt{\color{orangered}{1}}+2\color{orangered}{\cdot1}\]and x=1+h,\[\large\rm f(\color{orangered}{1+h})=\sqrt{\color{orangered}{(1+h)}}+2\color{orangered}{(1+h)}\] These are the two pieces that will get plugged into your numerator.

OpenStudy (canada907cat):

okay so how would you solve that per say since I plugged it all in and got 2 and Im not sure if thats correct

zepdrix (zepdrix):

Well, after you plug everything in,\[\large\rm f'(1)=\lim_{h\to0}\frac{\sqrt{1+h}+2(1+h)-(\sqrt1+2)}{h}\]The numerator should simplify a little bit, ya?\[\large\rm f'(1)=\lim_{h\to0}\frac{\sqrt{1+h}+2h-1}{h}\]

zepdrix (zepdrix):

You got 2? 0_o hmm

OpenStudy (canada907cat):

I didnt get the simplified version so I did something wrong

zepdrix (zepdrix):

There is a bunch more simplification, it's going to feel sort of weird.

OpenStudy (canada907cat):

I keep cancelling out the h's so it keeps equally 2 so i get sqrt1 +2-1 so Im clearly doing something wrong

zepdrix (zepdrix):

\[\large\rm f'(1)=\lim_{h\to0}\frac{\sqrt{1+h}+2h-1}{h}\]Direct substitution does not seem to work from this point, ya? We still have that pesky h in the bottom. Let's split the numerator into two groups,\[\large\rm f'(1)=\lim_{h\to0}\frac{\sqrt{1+h}+(2h-1)}{h}\]To get rid of that square root, we can apply conjugation: Multiplying the numerator and denominator by the conjugate of the numerator.

OpenStudy (canada907cat):

so 1+h+(2h-1)^2/h^2?

zepdrix (zepdrix):

No no. We're not multiplying like this,\[\large\rm f'(1)\ne\lim_{h\to0}\frac{\sqrt{1+h}+(2h-1)}{h}\color{royalblue}{\left(\frac{\sqrt{1+h}-(2h-1)}{h}\right)}\]we're multiplying like this,\[\large\rm f'(1)=\lim_{h\to0}\frac{\sqrt{1+h}+(2h-1)}{h}\color{royalblue}{\left(\frac{\sqrt{1+h}-(2h-1)}{\sqrt{1+h}-(2h-1)}\right)}\]

zepdrix (zepdrix):

So we shouldn't be getting the h^2 in the bottom, ya?

zepdrix (zepdrix):

And remember, multiplying conjugates gives you the `difference` of squares. So we should be getting subtraction between the split that we made.

OpenStudy (canada907cat):

oh you mean the entire thing oooooh okay i see where you were going with that. means that what would it look like after you multiply it since Im really lost at this point

zepdrix (zepdrix):

Here's a hint when applying this conjugate trick. `Only expand out the conjugates.` So in this case, it might be tempting to multiply the h in the denominator by that big mess, but don't! Just leave the multiplication sitting there. Don't expand the bottom.\[\large\rm f'(1)=\lim_{h\to0}\frac{(1+h)-(2h-1)^2}{h[\sqrt{1+h}-(2h-1)]}\]Your numerator was very close to being correct. You just had the wrong sign between them.

zepdrix (zepdrix):

Simplify the numerator, what do you get?

OpenStudy (canada907cat):

(1+h)-(2h-1)/h[sqrt1+h ] ?

zepdrix (zepdrix):

? I don't understand what you did..

zepdrix (zepdrix):

Oh oh I think I see what you're trying to do. No no, don't ever do that again :D lol

zepdrix (zepdrix):

\[\large\rm \frac{a+b}{c+b}\ne\frac{a+\cancel b}{c+\cancel b}\]bad bad bad. You can not do this.

OpenStudy (canada907cat):

I don't either honestly. You see I don't know what I am doing since my class is extremely vague so yea I apologise for now knowing. Also I kinda get why I cant do that but you see that how they taught us so I do it automatically

zepdrix (zepdrix):

Simplify the numerator,\[\large\rm (1+h)-(2h-1)^2=?\]

OpenStudy (canada907cat):

1+h-4h-4h^2-1?

OpenStudy (canada907cat):

it should be +1 sorry

zepdrix (zepdrix):

I don't think you distributed the negative properly. Or you didn't altogether? Hmm

zepdrix (zepdrix):

\[\large\rm (1+h)-(2h-1)^2\]expanding the square,\[\large\rm (1+h)-(4h^2-4h+1)\]

OpenStudy (canada907cat):

Oh geez I give up, I have another problem I have to do and that the same as this one and I still dont know what Im doing and oh yea I just combined it thats all

zepdrix (zepdrix):

Distributing the negative gives us,\[\large\rm (1+h)-4h^2+4h-1\]

OpenStudy (canada907cat):

oh okay I see. do we have to factor it?

zepdrix (zepdrix):

Combine like-terms,\[\large\rm -4h^2+5h\]That should be all that remains.

zepdrix (zepdrix):

In the numerator*

OpenStudy (canada907cat):

ah I see so would based on all of that, would the answer be 1/2?

zepdrix (zepdrix):

It's still too early to be guessing answers. Not sure why you're doing that :P Hmm

OpenStudy (canada907cat):

sorry its a multiple choice question so Im jsut guessing

zepdrix (zepdrix):

\[\large\rm f'(1)=\lim_{h\to0}\frac{(1+h)-(2h-1)^2}{h[\sqrt{1+h}-(2h-1)]}\]So our numerator expanded and simplified to give us,\[\large\rm f'(1)=\lim_{h\to0}\frac{4h^2-5h}{h[\sqrt{1+h}-(2h-1)]}\]We should probably pull an h out of each term in the numerator,\[\large\rm f'(1)=\lim_{h\to0}\frac{h(4h-5)}{h[\sqrt{1+h}-(2h-1)]}\]Ooo now we can get rid of that pesky h in the denominator! Do you see it?

zepdrix (zepdrix):

Oh sorry it's (4h+5) in the numerator.

OpenStudy (canada907cat):

okay yes I see that but how would we get rid of the rest? would we have to multiple it to remove the sqrt

zepdrix (zepdrix):

Well we have this nice cancellation happening,\[\large\rm f'(1)=\lim_{h\to0}\frac{\cancel h(4h+5)}{\cancel h[\sqrt{1+h}-(2h-1)]}\]In limit problems, any time you're able to mke a cancellation, go back to your first line of attack: Direct substitution. Something has changed, and we need to check and see if direct substitution will work from this point.

zepdrix (zepdrix):

\[\large\rm f'(1)=\lim_{h\to0}\frac{4h+5}{\sqrt{1+h}-(2h-1)}\]So what does direct substitution give you? Replacing h with 0.

OpenStudy (canada907cat):

5/1-(-1)=5/2?

zepdrix (zepdrix):

5/(1-(-1))=5/2? Yes, good job. That's our final answer.

OpenStudy (canada907cat):

ahhhh i see cool thank you for helping me with that now I have one more thats just like this one, but f(x)=sqrt(x+2) -3 and c=-1. Can you help me with this one as well? Its my last problem.

zepdrix (zepdrix):

This one will probably be a tad easier since we don't have multiple x's in the function.

OpenStudy (canada907cat):

oh yea thank you

zepdrix (zepdrix):

\[\large\rm f'(-1)=\lim_{h\to0}\frac{\color{royalblue}{\sqrt{x+h+2}-3}-\color{orangered}{(\sqrt{x+2}-3)}}{h}\]Here is our limit definition, with the two pieces plugged into the numerator.

zepdrix (zepdrix):

Oh we're evaluating this at x=-1, woops.

zepdrix (zepdrix):

\[\large\rm f'(-1)=\lim_{h\to0}\frac{\color{royalblue}{\sqrt{-1+h+2}-3}-\color{orangered}{(\sqrt{-1+2}-3)}}{h}\]

OpenStudy (canada907cat):

okay I have that all written down

OpenStudy (canada907cat):

so would we multiply it by the numerator to get rid of the sqrt?

zepdrix (zepdrix):

Let's simplify a little before moving too far. 2 and -1 gives us 1. Square root of 1 is 1.\[\large\rm f'(-1)=\lim_{h\to0}\frac{\color{royalblue}{\sqrt{1+h}-3}-\color{orangered}{(1-3)}}{h}\]We can simplify further still, ya?

OpenStudy (canada907cat):

sqrt1+5-1?

OpenStudy (canada907cat):

sorry I didnt meant to put the 5 in there

OpenStudy (canada907cat):

sqrt1+h -1

zepdrix (zepdrix):

sqrt(1+h)-1 ok good.

zepdrix (zepdrix):

That's really unclear when you don't write it with brackets :) lol

OpenStudy (canada907cat):

ah sorry ill try to do that

zepdrix (zepdrix):

\[\large\rm f'(-1)=\lim_{h\to0}\frac{\sqrt{1+h}-1}{h}\]Ok conjugation from this step seems like a good idea :)

OpenStudy (canada907cat):

so 1+h+1/h sqrt(1+h) -1

zepdrix (zepdrix):

Multiplying conjugates should give you the `difference` of squares. So we shouldn't have the second addition in the numerator. (1+h)-1 And what's going on in the denominator? Again, you need to learn to use brackets... h(sqrt(1+h)-1) But this is still a little off. We multiplied top and bottom by sqrt(1+h)+1), right?

OpenStudy (canada907cat):

ah right okay so let me write that out to visually it.

zepdrix (zepdrix):

\[\large\rm f'(-1)=\lim_{h\to0}\frac{\sqrt{1+h}-1}{h}\color{royalblue}{\left(\frac{\sqrt{1+h}+1}{\sqrt{1+h}+1}\right)}\]Ya, it uses a lot of paper to write it out, but I like to see that step when I'm working. :)

OpenStudy (canada907cat):

so (1+h)-1/h(sqrt(1+h)+1)

zepdrix (zepdrix):

Ok great. We can probably drop the brackets on the (1+h) since nothing is being applied to them,\[\large\rm f'(-1)=\lim_{h\to0}\frac{1+h-1}{h(\sqrt{1+h}+1)}\]and then simplify the numerator further.

OpenStudy (canada907cat):

so just h?

zepdrix (zepdrix):

Just h left up top? Ok good. Cancel some h's.

zepdrix (zepdrix):

\[\large\rm f'(-1)=\lim_{h\to0}\frac{\cancel h}{\cancel h(\sqrt{1+h}+1)}\]Gotta be careful when using this idea of "cancelling", think for a sec, what is left in the numerator?

OpenStudy (canada907cat):

1?

zepdrix (zepdrix):

Good good good. It's not zero.

zepdrix (zepdrix):

\[\large\rm f'(-1)=\lim_{h\to0}\frac{1}{\sqrt{1+h}+1}\]And once again, we can try direct substitution.

OpenStudy (canada907cat):

okay let me do that now

OpenStudy (canada907cat):

1/2?

zepdrix (zepdrix):

Yay good job

zepdrix (zepdrix):

cat cat kitty cat meow meow, practice those Algebra skills!

OpenStudy (canada907cat):

Lol thank you oh so much! I had about three mental breakdowns trying to do this and you just helped me get to the finish line. I really appreciate it.

zepdrix (zepdrix):

np take a break, drink a beer unless you're not old enough to drink XD then dont

OpenStudy (canada907cat):

Haha lol will do (except the drinking, I'm not quite there yet) and thank you and I appreciate the gesture.